OpenStudy (anonymous):

Daniel opened a car wash station in his neighborhood. He recorded the number of cars washed each week for the first six weeks. The data are shown as a scatter plot below. Based on the line of best fit, which of these would be the best prediction for the number of cars washed in Week 9? Answer 8 24 28 19 3 points Question 23

OpenStudy (zepp):

Where's the scatter plot?

OpenStudy (anonymous):

i did it do you see it

OpenStudy (zepp):

That arrow line would be your 'perfect line'; can you find the equation of this line?

OpenStudy (zepp):

There's two clear points: (0,0) and (3,8). Can you find the equation of this line by using these two points?

OpenStudy (anonymous):

no

OpenStudy (zepp):

Ok let's try to find the slope first by using the formula \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

using those two post

OpenStudy (zepp):

Yes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which order do i put first

OpenStudy (zepp):

Okay, I gave you the two points in the graph; (0,0) and (3,8) Now you can just name them and put them in the slope equation \(x_1=0;~y_1=0\) (Point (0,0)) \(x_2=3;~y_2=8\) (Point (3,8))

OpenStudy (anonymous):

so that would be -5,0

OpenStudy (zepp):

\[m=\frac{y_2-y_1}{x_2-x_1}=\frac{8-0}{3-0}=\frac{8}{3}\] You should get that, how did you arrive to -5,0)?

OpenStudy (anonymous):

i did 0-0 and 3-8

OpenStudy (zepp):

Look careful at the formula.

OpenStudy (zepp):

\(x_1\) is 0 \(x_2\) is 3

OpenStudy (anonymous):

oh so its 8/3

OpenStudy (zepp):

Yes.

OpenStudy (zepp):

Now by using a point and the slope we just found, find the equation of this line.

OpenStudy (anonymous):

idk its hard

OpenStudy (zepp):

Use the point-slope form \[(y-y_1)=m(x-x_1)\] Where \(\large (x_1,y_1)\) is a point on the line.

OpenStudy (anonymous):

idk

OpenStudy (zepp):

Here we found that the slope is \[\frac{8}{3}\] Let's use the points (0,0)\[(y-y_1)=m(x-x_1)\\y-0=\frac{8}{3}x\\y=\frac{8}{3}x\]

OpenStudy (zepp):

This would be our equation. Now they want the prediction for week 9

OpenStudy (anonymous):

yes

OpenStudy (zepp):